OpenStudy (anonymous):

how do i find the difference quotient using f(x) - f(a)/x-a? With x^3 at a=3

OpenStudy (solomonzelman):

your function is f(x)=x^3 and your a is 3, so tell me 2 things so that we can proceed. f(3)=? f(x+h)=?

OpenStudy (solomonzelman):

what you know tell me, what you don't, ask.

OpenStudy (anonymous):

well really i cant figure much out

OpenStudy (solomonzelman):

\(\large\color{slate}{ f(x)=x^3 }\), so if I wanted to find f(4), this is what i would do. \(\large\color{slate}{ f(4)=(4)^3 }\) \(\large\color{slate}{ f(4)=64 }\)

OpenStudy (anonymous):

f(3) = 27

OpenStudy (solomonzelman):

yes.

OpenStudy (anonymous):

(x+h)^3

OpenStudy (solomonzelman):

yes.

OpenStudy (solomonzelman):

\(\large\color{slate}{ f(x+h)=(x+h)^3 }\)

OpenStudy (solomonzelman):

So lets write the quotient.

OpenStudy (anonymous):

i cant get passed (x^3)-27/(x-3)

OpenStudy (anonymous):

the full x+h quotient?

OpenStudy (solomonzelman):

yes

OpenStudy (solomonzelman):

\(\large\color{slate}{ \displaystyle\frac{x^3-27}{x-3} }\)

OpenStudy (solomonzelman):

factor the top, can you do that?

OpenStudy (solomonzelman):

no, no x+h, that was my mistake, siorry and disregard the x+h

OpenStudy (solomonzelman):

\(\large\color{slate}{ \displaystyle\frac{x^3-27}{x-3} }\) is right

OpenStudy (anonymous):

no. that's where i'm stuck...

OpenStudy (anonymous):

i can't figure out how it factors...

OpenStudy (solomonzelman):

\(\large\color{slate}{ \displaystyle a^3-b^3=(a-b)(a^2+ab+b^2) }\)

OpenStudy (solomonzelman):

now, \(\large\color{slate}{ \displaystyle\frac{x^3-3^3}{x-3} }\) Apply \(\large\color{slate}{ \displaystyle a^3-b^3=(a-b)(a^2+ab+b^2) }\) to the top.

OpenStudy (anonymous):

so (x-3) (x^2 +3x +9)

OpenStudy (solomonzelman):

questions?

OpenStudy (anonymous):

well thats what i kept getting but it looked wrong for some reason.

OpenStudy (anonymous):

so the simplified answer comes out to be x^2 +3x +9?

OpenStudy (solomonzelman):

yup

OpenStudy (anonymous):

thank you very much :)

OpenStudy (anonymous):

i hope i remember that factoring rule later

OpenStudy (solomonzelman):

btw, if you are taking calculus, \(\large\color{slate}{ \displaystyle\lim_{x \rightarrow ~a}\frac{f(x)-f(a)}{x-a} }\) is a derivative of a function f(x) at x=a, or in other words f'(a).

OpenStudy (solomonzelman):

whether you are in calc or not, yw.....

OpenStudy (solomonzelman):

\(\large\color{slate}{ x^2 +3x +9 }\), but i think it isn;t over

OpenStudy (solomonzelman):

your x=3, right, so you have to plug that in.

OpenStudy (anonymous):

well a=3

OpenStudy (solomonzelman):

you couln't plug in 3 for x, becuase you would be getting a zero in the denominator, but now you can.

OpenStudy (solomonzelman):

yes, so \(\large\color{slate}{ x^2 +3x +9 }\) at x=3, means that \(\large\color{slate}{ (3)^2 +3(3) +9 }\)

OpenStudy (anonymous):

but it would just be 27 i dont think thats what she wants....

OpenStudy (solomonzelman):

it is not 27

OpenStudy (solomonzelman):

you are plugging 3 into the quotient, not just finding the f(3)

OpenStudy (solomonzelman):

we need to evaluate \(\large\color{slate}{\displaystyle\frac{x^3-27}{x-3}}\) at x=3. We couldn't do that, because as we plug in 3 we get 0 for the denominator. So we went ahead and factored the numerator, and got: \(\large\color{slate}{\displaystyle\frac{(x-3)(x^2 +3x +9 )}{x-3}}\) now x-3 cancels out. and you remain with \(\large\color{slate}{\displaystyle x^2 +3x +9 }\)

OpenStudy (solomonzelman):

and \(\large\color{slate}{\displaystyle x^2 +3x +9 }\) can surely be evaluated at x=3

OpenStudy (solomonzelman):

any confusion ?

OpenStudy (anonymous):

which is 27...?

OpenStudy (anonymous):

9 + 9 + 9 = 27

OpenStudy (solomonzelman):

oh, yes, right

OpenStudy (solomonzelman):

very good.

OpenStudy (solomonzelman):

27 is indeed the cprrect ans.

OpenStudy (solomonzelman):

don't panic if it is too simple, lol. It is a good thing not a bad thing.